OpenStudy (anonymous):

how do you solve this system of equations x=2y-8 4x+y=13

OpenStudy (anonymous):

x=2 and y=5

OpenStudy (anonymous):

how did you get that CAN YOU DRAW IT! pleassseee

OpenStudy (anonymous):

2y=x+8 y=13-4x(*-2) -2y=-26+8x So, 2y=x+8 -2y=-26+8x 0=9x-18 9x=18 x=2 2y=x+8 2y=(2)+8 2y=10 y=5

OpenStudy (anonymous):

Tried, and failed utterly, to do a drawing, but here's my way of going about the problem. #1: x = 2y - 8 #2: 4x + y = 13 Now, the trick is to express x in a succinct manner. You start my manipulating the second equation: 4x + y = 13 4x = 13 - y x = (13 - y) / 4 Now that we have successfully expressed x in terms of y, we proceed to IMPLEMENT this new expression into #1: x = 2y - 8 (13 - y) / 4 = 2y - 8 13 - y = 4(2y - 8) 13 - y = 8y - 32 13 = 9y - 32 9y - 32 = 13 9y = 45 y = 5 With me so far? :) Good! Now that we have successfully found the value of y, we need to find the value of x. To do this, we simply apply our newly found value of y into #1: x = 2y - 8 x = 2 * 5 - 8 x = 10 - 8 x = 2 And that is how it's done! :)

OpenStudy (anonymous):

thanks! (:

OpenStudy (anonymous):

yeah, I tried to be very simples buttt :]